OpenStudy (anonymous):

1/4 of a candy bar is eaten, then 1/3 of the candy bar is eaten, what fraction of the candy bar is left?

hero (hero):

Hint: \[(1 - \frac{1}{4}) - \frac{1}{3}\]

hero (hero):

2nd Hint: \[(\frac{4}{4} - \frac{1}{4}) - \frac{1}{3}\]